# The BPgWSP test: a Bayesian Weibull Shape Parameter signal detection test for adverse drug reactions

### v1

@@ Abstract
We develop a Bayesian Power generalized Weibull shape parameter (PgWSP) test as statistical method for signal detection of possible drug-adverse event associations using electronic health records for pharmacovigilance. The Bayesian approach allows the incorporation of prior knowledge about the likely time of occurrence along time-to-event data. The test is based on the shape parameters of the Power generalized Weibull (PgW) distribution. When both shape parameters are equal to one, the PgW distribution reduces to an exponential distribution, i.e. a constant hazard function. This is interpreted as no temporal association between drug and adverse event. The Bayesian PgWSP test involves comparing a region of practical equivalence (ROPE) around one reflecting the null hypothesis with estimated credibility intervals reflecting the posterior means of the shape parameters. The decision to raise a signal is based on the outcomes of the ROPE test and the selected combination rule for these outcomes.
The development of the test requires a simulation study for tuning of the ROPE and credibility intervals to optimize specificity and sensitivity of the test. Samples are generated under various conditions, including differences in sample size, prevalence of adverse drug reactions (ADRs), and the proportion of adverse events. We explore prior assumptions reflecting the belief in the presence or absence of ADRs at different points in the observation period. Various types of ROPE, credibility intervals, and combination rules are assessed and optimal tuning parameters are identified based on the area under the curve. The tuned Bayesian PgWSP test is illustrated in a case study in which the time-dependent correlation between the intake of bisphosphonates and four adverse events is investigated.
@@ 1 Introduction
Making assumptions about the generating process of an outcome in the development of a test of association can improve the performance of this test in terms of power, specificity, and sensitivity by enabling the use of more statistical information included in the data. In the field of pharmacovigilance (surveillance of occurrence of adverse events (AEs) related to prescribed drugs (Meyboom et al.,, 1997), this approach has been used by Cornelius and colleagues (Cornelius et al.,, 2012; Sauzet and Cornelius,, 2022) to develop a family of signal detection tests (test of association between the occurrence of an AE and a prescribed drug) which is based on an assumption about time dependence on drug prescription. More specific, Cornelius et al., (2012) proposed tests aiming for detection of a non-constant hazard function by modeling the time to occurrence of event from the time of prescription to flag a possible drug-event association. However, this is not the only type of information about a generating process that can be incorporated into a test. Prior knowledge regarding the distribution parameters of the generating process (e.g. mean time of occurrence of an event) can be incorporated in the development of a test using a Bayesian approach. The aim of this article is to combine both approaches by developing a Bayesian test to detect non-constant hazard functions for a time-dependent process incorporating prior knowledge on the distribution of the time to the occurrence of events. While the principle is rather general, we present this test in the context of pharmacovigilance which we briefly describe below, and review some existing methods of signal detection that assume a temporal relationship or use a Bayesian approach.
After release of a drug on the market, pharmacovigilance monitors the occurrences and changes of known adverse drug reactions (ADRs) as well as aims to detect new ADRs in the population. An ADR is an unintended, harmful response to a medicine (EMA,, 2024).
The investigation of a harm profile of a drug consists of multiple steps: screening the range of AEs that occur along prescribed drugs, performing a signal detection test to investigate the association for a drug-AE pair and, in case a signal is raised, deeper investigation of any causal relationship between the two which involves a complex epidemiological analysis (Meyboom et al.,, 1997).
Longitudinal electronic health records (EHRs) gathered by health insurance companies and health regulatory organizations such as the Sentinel system established by the FDA (Food and Drug Administration,, 2022) provide regularly collected longitudinal data about diagnoses, treatments, and medication prescribed by physicians, as well as AEs (Suling and Pigeot,, 2012; Sturkenboom and Schink,, 2021).
@@
With the collection of EHR data, statistical methods to exploit them have emerged. These include among others extensions of disproportionality approaches comparing rates of occurrence between a single drug and the rest of the dataset (Zorych et al.,, 2013) for application to longitudinal data (Schuemie,, 2011), temporal pattern analysis (Norén et al.,, 2010), exposure models (Dijkstra et al.,, 2024) and more (see for example the review by Coste et al., (2023)).
As mentioned above Cornelius et al., (2012) and Sauzet and Cornelius, (2022) considered the temporal dynamics of the risk to experience a specific AE in their family of Weibull shape parameter (WSP) family of tests. For this a Weibull type distribution parametric hazard function is fitted to the data. A signal is raised following the rejection of a null-hypothesis assuming a constant hazard which means that a sufficiently large change in risk over time can be interpreted as a signal. Unlike disproportionality methods, no group comparisons are performed.
@@
Bayesian statistical methods for signal detection have also been developed. But methods like the Empirical Bayes Gamma Mixture/ Gamma-Poisson shrinker or the Bayesian confidence propagation neural network are Bayesian extensions of disproportionality measures used to obtain signals from spontaneous reporting databases (Bansal et al.,, 2024) which do not consider the time of the event occurrence.
@@
The objective of this work is to combine an assumption of temporal process with the incorporation of prior knowledge about distribution parameters of this process in the development of Bayesian Weibull shape parameter tests. To do so, we will generalize the proposed power generalized Weibull shape parameter (PgWSP) test by Sauzet and Cornelius, (2022) into a Bayesian framework to develop and tune an HDI+ROPE type test of signal detection.
The remaining article is structured as follows. In section 2 the theoretical concept of the Bayesian PgWSP (BPgWSP) test is presented. In section 3, we conduct a simulation study to tune the parameters defining the BPgWSP test in a chosen framework of simulation scenarios. Finally, in section 4, we apply our tuned BPgWSP test to data on women prescribed with bisphosphonates from the The Health Improvement Network (THIN) (THIN,, 2024) database in a case study.
@@ 2 Methodological concept for the Bayesian signal detection test
The parametric distribution is defined by a parameter vector $\Theta$. We make use of a subset of parameters $\mathcal{N}\subset\Theta$ on which to define the null-hypothesis of constant hazard and develop the Bayesian signal detection test based on an updated distribution for $\mathcal{N}$ using the HDI+ROPE approach (see sec. 2.2). For this we will perform the three steps below:
@@
    describe the BPgWSP test using the HDI+ROPE approach,
@@
2.1 Power generalized Weibull distribution:
The survival function of the PgW distribution is given by:
@@ 2.1 Power generalized Weibull distribution:
The hazard function is given by:
$h(t)=\frac{\nu}{\gamma\theta^{\nu}}\cdot t^{\nu-1}\cdot\left[1+\left(\frac{t}{\theta}\right)^{\nu}\right]^{\frac{1}{\gamma}-1},$ (2)
(see Bagdonavicius and Nikulin, (2001, p.8); Nikulin et al., (2016, p.8)). The PgW distribution has three parameters $\Theta_{PgW}=\{\theta,\nu,\gamma\}$ two of which are shape parameters such that $\mathcal{N}_{PgW}=\{\nu,\gamma\}$.
@@
The two shape parameters allow for a range of hazard function forms including constant, monotoneously in- or decreasing and unimodal or bathtub shaped (see Figure 2).
A constant hazard is given if and only if both shape parameters are equal to one so that the null-hypothesis of constant hazard can be formulated as:
@@ 2.2 Development of Bayesian signal detection test using the HDI+ROPE approach
Prior and posterior distribution
The prior distribution of $\Theta$ should reflect whether the AE of interest is suspected to be an ADR (non-constant hazard) and if so, when it is expected to occur (specifying $\mathbf{E}[t]$) a priori by medical experts. This can be translated into the corresponding expected hazard form (comp. Figure 2) parameterized in terms of prior mean values for the parameters in $\Theta$. The prior distribution of the scale parameter reflects the AE background rate. The priors of the shape parameters $\mathcal{N}\subset\Theta$ determine the constancy or non-constancy and general trend of the hazard over time.
Based on the specified prior means (and a chosen standard deviation) as well as a distributional choice (eg. a Lognormal distribution for each parameter in $\Theta$), the prior distribution $p_{0}(\Theta)$ will be updated with the data available. For time-to-event data the posterior distribution $p_{1}(\Theta|t)$ is proportional to
@@
with $\mathcal{L}(t|\Theta)$ being the likelihood function in the time-to-event framework such that all subjects $i$ not having an AE during the observation time (denoted by $d_{i}=0$) are considered in terms of the survival function $S(t_{i})$ and all subjects $i$ experiencing an AE (denoted by $d_{i}=1)$ are taken into account in terms of the density function $f(t_{i})=S(t_{i})\cdot h(t_{i})$.
HDI+ROPE method
Based on the posterior distribution of $\mathcal{N}$, the HDI+ROPE test (Kruschke,, 2015, 2018) applied to each shape parameter and combination of all single shape parameter test results provide two possible outcomes: either we lean towards the null-hypothesis of constant hazard or the alternative hypothesis of a non-constant hazard. The test setup depends on
  • the region of practical equivalence (ROPE) to the null value, in our case one, typically defined as an interval around the null value and
  • a credibility interval (CI) based on posterior distribution parameters in $\mathcal{N}$, here either a highest density interval (HDI) as proposed by Kruschke, (2015) or an equal tailed interval (ETI). See sec. 3.3 for the formal definition of the ETI and HDI.
Figure 3: HDI+ROPE decision rules based on graphic from Kruschke, (2018, Fig. 1).
In our setting, the HDI+ROPE test is applied to each parameter $\nu\in\mathcal{N}$ and provides one of the following interim results (compare to Figure 3):
@@
    is accepted,
  • if ROPE and CI do not intersect at all, then the single parameter null-hypothesis is rejected,
  • if ROPE and CI partly overlap, then no decision is made (undecided).
Given the single parameter test outputs, the combined test result is concluded according to a combination rule, for instance the condition directly implied by Equation 3:
@@
To complete the test development, we need to make decisions about the distributional choice for the prior, the position and width of the ROPE and the CI of the BPgWSP test and the combination rule. This will be done in the following section.
@@ 3 Simulation study for test development and tuning
A method to finalize the BPgWSP test theoretically described in section 2 is to perform a simulation study. By applying a range of prior distributions, choices for position and width of the ROPE, position and width of the CI and the combination rules, we make decisions about the optimal tuning parameters for the BPgWSP test such that a performance parameter based on sensitivity and specificity of the test is maximal across a range of sample scenarios.
@@ 3.1 Data generating process
The set of simulation scenarios is determined by the
  • censoring time (365 days) which sets the end of the observation period,
@@
  • expected event-time of the ADR if the ADR rate is above zero (1st, 2nd, 3rd quarter of observation period),
@@
The background observations are sampled using a Binomial distribution with a probability equal to the specified background rate
@@
Corresponding event-times are sampled using a Uniform distribution over the study period ranging from day one to day $365$.
@@
The number of events induced by the drug intake is sampled from a Binomial distribution with a probability equal to the background rate times the ADR rate.
@@
The event-times for the ADR cases are generated from a Normal distribution with an expected event-time $\mathrm{E}[t]$ at the first quarter, half, and third quarter of the study period and a standard deviation of $0.05$ times the censoring time at 365 days.
$t^{ADR}$ $\sim N(\mathrm{E}[t],0.05\cdot 365)$ (8)
@@
The sample is filled up to size $N$ with censored observations. They have the AE status zero and time point $365$ equaling the end of the study period.
3.2 Prior specifications for the Bayesian model
@@
Figure 4: Hazard functions obtained under the prior belief that the expected event-time is at the first, second, or third quarter of the observation period (left) or that the AE of interest is no ADR (right).
prior belief about prior means
expected event-time $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$
none $(1,1,1)$
1st quarter of OP (91 days) $(1,0.207,1)$
2nd quarter of OP (183 days) $(20,5.5,14)$
3rd quarter of OP (274 days) $(300,4,1)$
Table 1: Mean of prior distribution under implemented prior beliefs about expected event-time.
@@
Four different prior beliefs are implemented. An ADR is expected to occur with higher risk either around the first, second or third quarter of the observation period or not expected at all. These assumptions are represented by hazard functions of the distribution shown in Figure 4 based on chosen prior mean values for the scale $\theta$, and shapes $\nu$ and $\gamma$ listed in Table 1. The standard deviation of the parameters’ prior distributions is set to 10.
scale $\theta$ shape $\nu$ powershape $\gamma$ abbreviation
fixed to prior mean Gamma Gamma fix-gam-gam
Gamma Gamma Gamma gam-gam-gam
fixed to prior mean Lognormal Lognormal fix-log-log
Lognormal Lognormal Lognormal log-log-log
Table 2: Prior distributional assumptions for the scale $\theta$, and shapes $\nu$ and $\gamma$ of the PgW distribution.
For practical reasons, scale and shape parameters are assumed to be independent. By definition all parameters of the PgW distribution are positive. Four prior distributional choices matching these conditions are listed in Table 2. All parameters’ univariate priors follow either the Gamma or Lognormal distribution. In two of the four prior distribution choices, the scale parameter of the posterior remains equal to the prior (i.e. will not be updated).
For instance, the PgW model with log-log-log prior is be specified as follows:
tsimilar-$t\sim$ $PgW(\theta,\nu,\gamma)$ (9)
$\theta_{0}\sim Lognormal(\mu_{0}^{\theta},\sigma_{0}^{\theta})$ (10)
$\nu_{0}\sim Lognormal(\mu_{0}^{\nu},\sigma_{0}^{\nu})$ (11)
$\gamma_{0}\sim Lognormal(\mu_{0}^{\gamma},\sigma_{0}^{\gamma})$ (12)
where the prior meta parameters of the Lognormal distribution are the logarithm of the location $\mu$ and the logarithm of the scale $\sigma$ which can be calculated from the prior mean (specified in Table 1 and standard deviation (10) as
$\mu_{0}^{\theta}$ $=\log(\bar{\theta}_{0})-\log\left(\frac{sd(\theta_{0})^{2}}{\bar{\theta}_{0}^{2}}+1\right)$ (13)
$\sigma_{0}^{\theta}$ $=\sqrt{\log(\frac{sd(\theta_{0})^{2}}{\bar{\theta}_{0}^{2}}+1)}$ (14)
for $\theta$ and analogously for $\nu$ and $\gamma$.
A posterior estimate is obtained by Markov chain Monte Carlo (MCMC) sampling. We draw a posterior sample of size 40000 obtained from four chains of length 10000 after a burn-in phase of 1000 draws each.
@@
Each combination of sample scenario and prior specification leads to a total of 336 different simulation settings – 288 with a positive ADR rate and 48 controls to match in with only background events for the simulation study. Each simulation scenario’s data generation and posterior estimation is repeated 100 times.

### v3

@@ Abstract
We develop the Bayesian Power generalized Weibull shape parameter (BPgWSP) test as statistical method for signal detection of possible drug-adverse event associations using electronic health records for pharmacovigilance. The Bayesian approach allows the incorporation of prior knowledge about the likely time of occurrence along time-to-event data. The test is based on the shape parameters of the Power generalized Weibull (PgW) distribution. When both shape parameters are equal to one, the PgW distribution reduces to an exponential distribution, yielding a constant hazard function. This is interpreted as no temporal association between drug and adverse event (AE). The BPgWSP test involves comparing a region of practical equivalence (ROPE) around one reflecting the null hypothesis with estimated credibility intervals (CI) reflecting the posterior means of the shape parameters. The decision to raise a signal is based on the CI+ROPE tests and the selected combination rule for these outcomes.
The test development requires a simulation study for tuning of the ROPE and CIs to optimize specificity and sensitivity of the test. Samples are generated under various conditions, including differences in sample size, prevalence of adverse drug reactions (ADRs), and the proportion of AEs. We explore prior assumptions reflecting the belief in the presence or absence of ADRs at different points in the observation period. Various types of ROPE, CIs, and combination rules are assessed, and optimal tuning parameters are identified based on the area under the curve. The tuned BPgWSP test is illustrated in a case study in which the time-dependent correlation between the intake of bisphosphonates and four AEs is investigated.
Julia Dyck [^1] Department of Business Administration and Economics, Bielefeld University, Universitätsstraße 25, Bielefeld, 33615, Germany.***Corresponding author: j.dyck@uni-bielefeld.de and Odile Sauzet11footnotemark: 1222School of Public Health, Bielefeld University, Universitätsstraße 25, Bielefeld, 33615, Germany.
@@ 1 Introduction
In the field of pharmacovigilance, signal detection methods are used to identify new potential adverse drug reactions1 (ADRs) among adverse events 2 (AEs) occurring in the population by testing for an association between the occurrence of an AE and a prescribed drug.3 Signal detection is important to keep a drug‘s harm profile updated and can result in adjustments of the prescription labelling or even a recall of the product from the market. 4, 5, 6
@@
With the collection of longitudinal electronic health records7, 8 (EHRs) (regularly collected data about diagnoses, treatments, and medication prescribed by physicians, as well as AEs), statistical methods to exploit them have emerged. These include disproportionality approaches adjusted for longitudinal data comparing rates of occurrence between a single drug and the rest of the dataset 9, 10, temporal pattern analysis 5, exposure models 11 and more (see for example the review by Coste et al. 12).
Sauzet and Cornelius13, 6 developed a family of signal detection tests based on an assumption about time dependence on the drug prescription. More specific, their tests aim for detection of a non-constant hazard function (depicting the risk over time) by modelling right-censored time-to-event data starting at drug prescription to flag a possible drug-event association. Making assumptions about the generating process of an outcome (more specifically the hazard function) in the development of a test of association can improve the performance of this test in terms of power, specificity, and sensitivity by enabling the use of more statistical information included in the data. However, this is not the only type of information that can be incorporated into a test.
@@
Prior knowledge regarding the distribution parameters of the generating process (e.g. mean time of occurrence of an event) might be incorporated into the development of a test using a Bayesian approach. Bayesian statistical methods like the Empirical Bayes Gamma Mixture, also called Gamma-Poisson shrinker, or the Bayesian confidence propagation neural network are extensions of disproportionality measures used to obtain signals from spontaneous reporting databases.14, 15 These, in turn, do not consider the time of the event occurrence.
@@
This article aims to combine the assumption of a temporal process and the incorporation of prior knowledge about distribution parameters of this process. To do so, we will generalize the power generalized Weibull shape parameter (pWSP) test proposed by Sauzet and Cornelius 6 into a Bayesian framework to develop and tune an HDI+ROPE16 type test (a Bayesian equivalent of a hypothesis test) of signal detection which we call the Bayesian Power generalized Weibull shape parameter (BPgWSP) test.
@@ 2 Methodological concept for the Bayesian signal detection test
The method is conceptualized for right-censored time-to-event data. Each observation $i$ consists of the time $t_{i}\in(0,c)$ with $c$ being the censoring time and the binary status $d_{i}\in\{0,1\}$ indicating whether the AE occurred at $t_{i}$. The parametric distribution is defined by a parameter vector $\Theta$. We define the null-hypothesis of constant hazard based on a subset of parameters $\mathcal{N}\subset\Theta$ and develop the Bayesian signal detection test (see Figure 1) based on an updated distribution for $\mathcal{N}$ using the HDI+ROPE approach (see sec. 2.2). For development, we will perform the three steps below:
@@
    describe the BPgWSP test using the HDI+ROPE approach based on a Bayesian estimation,
@@
2.1 Power generalized Weibull distribution
The survival function of the PgW distribution17, 18 is
@@ 2.1 Power generalized Weibull distribution
The hazard function17, 18 is
$h(t)=\frac{\nu}{\gamma\theta^{\nu}}\cdot t^{\nu-1}\cdot\left[1+\left(\frac{t}{\theta}\right)^{\nu}\right]^{\frac{1}{\gamma}-1}.$ (2)
The PgW distribution has three parameters $\Theta_{PgW}=\{\theta,\nu,\gamma\}$ - two of which are shape parameters such that $\mathcal{N}_{PgW}=\{\nu,\gamma\}$.
@@
The two shape parameters allow for a range of hazard function forms including constant, monotoneously in- or decreasing, and unimodal or bathtub shaped (see Figure 2).
A constant hazard is given if and only if both shape parameters are equal to one, so that the null-hypothesis of constant hazard can be formulated as
@@ 2.2 Development of Bayesian signal detection test using the HDI+ROPE approach
The prior distribution of $\Theta$ should reflect whether the AE of interest is suspected to be an ADR (by choosing a parameter combination yielding a non-constant hazard) and if so, when it is expected to occur a priori by medical experts (by specifying parameters $\Theta$ that lead to the a priori intended $\mathbf{E}[t]$). The scale parameter reflects the AE background rate. The shape parameters $\mathcal{N}\subset\Theta$ determine the constancy or non-constancy and general trend of the hazard over time.
For practicality, univariate prior distributions for each parameter out of $\Theta$ are specified. Concerning implementation, not $\mathbf{E}[t]$ but the prior means of all parameters are set, followed by checking whether the hazard form and $\mathbf{E}[t]$ resulting from the chosen prior means reflect the prior belief to a sufficient extent in a trial-and-error manner.
Based on the specified prior means (and a chosen standard deviation) as well as a distributional choice (e.g. the product of univariate lognormal distributions for each parameter in $\Theta$), the prior distribution $p_{0}(\Theta)$ will be updated yielding a posterior distribution. For time-to-event data the posterior distribution $p_{1}(\Theta|t)$ is proportional to
@@
being the likelihood function in the time-to-event framework such that all subjects $i$ not having an AE during the observation time (denoted by $d_{i}=0$) are considered in terms of the survival function $S(t_{i})$ and all $i$ experiencing an AE (denoted by $d_{i}=1)$ are taken into account in terms of the density function $f(t_{i})=S(t_{i})\cdot h(t_{i})$.
Based on the posterior distribution of $\mathcal{N}$, the HDI+ROPE test 19, 16 is applied to each shape parameter and combined, leading to two possible outcomes: either we lean towards the null-hypothesis of constant hazard or the alternative hypothesis of a non-constant hazard. The test setup depends on
  • the region of practical equivalence (ROPE) to the null value (in our case, one), typically defined as an interval around the null value and
  • a credibility interval (CI) based on posterior distribution parameters in $\mathcal{N}$, here either a highest density interval (HDI) as proposed by Kruschke19 or an equal-tailed interval (ETI). See sec. 3.3 for the formal definition of the ETI and HDI.
Figure 3: HDI+ROPE decision rules based on graphic from Kruschke, Fig. 1.16
In our setting, the HDI+ROPE test is applied to each parameter $\nu\in\mathcal{N}$ and provides one of the following interim results:
@@
    is accepted (Figure 3b-c),
  • if ROPE and CI do not intersect at all, then the single parameter null-hypothesis is rejected (Figure 3a),
  • if ROPE and CI partly overlap, no decision is made (Figure 3d-f).
Given the single parameter test outputs, the combined test result is concluded according to a combination rule, for instance, the condition implied by Equation 3:
@@
To complete the test development, we need to make decisions about the prior distributional choice, the position and width of the ROPE and the CI of the BPgWSP test and the combination rule.
@@ 3 Simulation study for test development and tuning
A method to finalize the BPgWSP test, theoretically described in sec. 2, is to perform a simulation study. By applying a range of prior distributions, choices for position and width of the ROPE, position and width of the CI, and combination rules, we make decisions about the optimal tuning parameters for the BPgWSP test such that a performance parameter based on sensitivity and specificity of the test is maximal across a range of sample scenarios. The frequentist pWSP test approach developed by Sauzet and Cornelius 6 with credibility levels to be tuned is considered in the simulations for comparison reasons.
@@ 3.1 Data generating process
The sample scenarios are determined by the
  • censoring time (365 days) setting the end of the observation period (OP),
@@
  • expected event-time of the ADR if the ADR rate is above zero (1st, 2nd, 3rd quarter of OP),
@@
The background observations are sampled from a binomial distribution with a probability equal to the specified background rate
@@
Corresponding event-times are sampled from a uniform distribution over the OP ranging from day one to $365$.
@@
The number of events induced by the drug is sampled from a binomial distribution with the probability being the background rate times ADR rate.
@@
Event-times for the ADR cases are generated from a normal distribution with an expected event-time $\mathrm{E}[t]$ at the first, second, and third quarter of the OP and a standard deviation of $0.05$ times the censoring time at 365 days.
$t^{ADR}$ $\sim\mathcal{N}(\mathrm{E}[t],(0.05\cdot 365)^{2})\$(8)
@@
The sample is filled up to size $N$ with censored observations. They have the AE status zero and time point $365$ (end of the OP).
3.2 Prior specifications and model estimation
@@
Figure 4: Hazard functions obtained under the prior belief that the expected event-time is at the first, second, or third quarter of the OP (left) or that the AE of interest is no ADR (right).
@@
Four different prior beliefs are implemented. An ADR is expected to occur with higher risk either around the first, second, or third quarter of the OP or not expected at all. These assumptions are represented by hazard functions of the distribution shown in Figure 4 based on chosen prior mean values for the scale $\theta$, and shapes $\nu$ and $\gamma$ listed in Table 1. The standard deviation of all parameters is set to 10. See the supplementary material for details on the implementation.
prior belief about a priori prior means
expected event-time $E[t]$ $(\bar{\theta}_{0},\bar{\nu}_{0},\bar{\gamma}_{0})$
none - $(1,1,1)$
1st quarter of OP 91 days $(1,0.207,1)$
2nd quarter of OP 183 days $(20,5.5,14)$
3rd quarter of OP 274 days $(300,4,1)$
Table 1: Mean of prior distribution under implemented prior beliefs about expected event-time. Prior standard deviation is set to 10 for all parameters in all prior belief cases.
To obtain the posterior estimate the No-U-Turn sampler is used 20. We draw a posterior sample of size 40 00040\,000 obtained from four chains of length 10 00010\,000 after a burn-in phase of 1 0001\,000 draws per estimation. In practice, model diagnostics should be used to check convergence and sampling efficiency before using it for testing.21 This is not done within the simulations due to the automated procedure, but showcased in the supplementary material.
@@
Each combination of sample scenario and prior specification leads to a total of 168 simulation settings – 144 with a positive ADR rate and 24 controls with only background events to match. Each simulation scenario’s data generation and posterior estimation is repeated 100 times.
The pWSP test does not require a prior specification, such that the number of simulation scenarios is reduced to 21 – 18 scenarios with a positive ADR rate and three controls. Data generation and maximum likelihood estimation are performed 800 times each.